# A modular perspective to the jet suppression from a small to large radius in very high transverse momentum jets

### v4

@@ I Introduction
The extremely hot and dense conditions created at the start of the Big Bang led to what we now understand as the soup of deconfined state of the partons, the quark-gluon plasma (QGP) [1, 2, 3]. The QGP is of such great interest that, to study the properties of this state of matter, experimentalists have spent decades building advanced detectors capable of recreating this extremely dense, soupy state. The Relativistic Heavy-Ion Collider (RHIC) [4, 5] and the Large Hadron Collider (LHC) [6, 7, 8, 9] conduct heavy-ion collisions where the QGP is created for very short periods. During this time, the QGP medium has a substantial impact on the propagation and modification of the parton shower. In particular, the high-${p_{T}}$ jets produced in these heavy ion collisions undergo strong yield suppression and medium modification which are together referred to as jet quenching phenomena [10, 11, 12]. Jets modification has been studied in nucleus-nucleus collisions relative to proton-proton collisions to probe the properties of the QGP via constraints from model-to-data agreement  [13, 14, 15, 16]. The measurement of the nuclear modification factor for jets as well as inclusive jet fragmentation functions has revealed many important characteristics of the quark-gluon plasma [17, 18, 19, 20]. It provides strong confirmation of the interaction of partons with the deconfined plasma, the resulting medium modifications, and the eventual transition to hydrodynamic behavior in the medium [8]. Since there are many conclusive studies based on jet-${R_{AA}}$ [21, 22, 23, 24, 25], our effort here is to push the limits of the current event generators and various energy loss models for a better description of jet quenching phenomena at very high transverse momenta and broader jet cones with the multi-stage evolution of the parton shower using the Jet Energy-loss Tomography with a Statistically and Computationally Advanced Program Envelope (JETSCAPE) framework (version 3.6.5)  [26].
@@ II Results
This paper covers the collision energy of ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for the most central (0-10%) Pb-Pb collisions and includes a comparison with selected experimental data from the ATLAS and the CMS collaborations for high-${p_{T}}$ jets. The energy loss depicted in this work is achieved by the coupling of MATTER with LBT model. The secondary LBT model (which handles the low virtuality phase) remains the same throughout until and unless specified. Fig. 1 (top panel) shows the ${p+p}$ collision results for inclusive jet spectra at ${\sqrt{s}=5.02~{}\rm TeV}$ for ${|y_{jet}|<2.8}$, which follow the JETSCAPE PP-19 tune [52], are then compared to the experimental data from the ATLAS [24]. In this study, we have set running coupling constant ($\alpha_{\mathrm{s}}$) to 0.35 for all the calculations.
@@
These results strongly support the fact that the MUSIC model is not only adequate but also highly effective in describing the experimental observations even at higher jet-${p_{T}}$ in the most central collisions. In contrast to the VISHNU model [54], MUSIC’s hydrodynamic evolution allows for a more accurate representation of the medium’s longitudinal expansion and viscous effects. This leads to a recurring trend of consistent predictions, indicating that MUSIC outperforms other hydrodynamic models in capturing the intricate dynamics of jet interactions within the quark-gluon plasma, especially at high energy scales reaching up to 1 TeV. The MUSIC model shows low sensitivity to coherence effects that dominate at high transverse momentum, compared to VISHNU (2+1)D. This enhances the reliability of MUSIC model for analyzing the jet-suppresions phenomenon at high-${p_{T}}$.
@@ II.1 Exploring the MARTINI and AdS/CFT models
Simulations of Pb-Pb collisions are carried out using the MATTER model for both vacuum and medium showers, integrated with three distinct time-ordered parton shower models: LBT, MARTINI and AdS/CFT. The transition between different energy loss models is performed independently for each parton, with the switching parameter $Q_{\mathrm{SW}}$ set at 2 GeV. Both the MARTINI [32] and the AdS/CFT [33] models are developed to handle the low virtuality phase and simulate the energy loss once the parton is transferred from the MATTER. We replace the successful LBT model with the MARTINI, to carry out the simulations for the same conditions [49] and compare the jet-${R_{AA}}$ with the experimental data from the CMS [35]. Similarly, the calculations are done by replacing the LBT with AdS/CFT model and compared with the experimental data.
@@ II.2 Jet radius (${R}$) and jet-${p_{\mathrm{T}}}$ dependence of the ${R_{AA}}$
A study by the CMS collaboration [35] enlightens the rigorous comparisons of predictions from quenched jet event generators like HYDJET++ [56], PYQUEN [57] and theoretical models used to replicate relativistic heavy-ion collisions to the experimental data for the jet-${R_{\mathrm{AA}}}$. The article concludes with the impression that although most state-of-the-art models have progressed, there is still considerable uncertainty regarding large-area jets.
The large jet radius also implies that the jet retains a significant proportion of the extensively distributed momentum and energy deposited in the plasma. Since the JETSCAPE framework has pioneered the realization of a multi-stage approach for modular-based energy loss, the motivation here is to challenge the JETSCAPE model that has so far met our expectations in describing the variety of data observed and to investigate the limitations of the current models.
@@
For ${\mathrm{300~{}GeV}\leq p_{T}^{jet}\leq\mathrm{400~{}GeV}}$ interval, Fig. 5 (left panel) shows the predictions by the different combination of energy loss models. However, we note that there is only one measurement done by the CMS [35] in this ${p_{T}}$ interval, i.e., for $R$ = 0.3. We observe a consistent and nearly identical trend across all three models, i.e., (MATTER + LBT), (MATTER + AdS/CFT), and (MATTER + MARTINI), each predicting a minimal enhancement in the double ratio and showing good agreement with the data. This also indicates that the energy loss of the partons throughout the various jet cone sizes is comparable among LBT, MARTINI, and AdS/CFT.
@@
In the extreme high $p_{T}^{jet}$ region, specifically for ${\mathrm{500~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$, shown in Fig. 5 (right panel), the experimental data reveal that double ratio is less than unity for intermediate R. This may indicate that highly boosted jets, which possess more intricate jet substructures, undergo marginally more suppression. However, all the models moderately overestimate the jet-${R_{\mathrm{AA}}}$ double ratio in this high $p_{T}^{jet}$ region. Among them, the (MATTER + AdS/CFT) model provides the closest agreement, deviating by only up to ${10\%}$, and also harmonises with the data within the uncertainties. In contrast, (MATTER + MARTINI) over-predicts the ratio although it follows the same general pattern observed in other $p_{T}^{jet}$ regions.
All the models predict that the double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) increases with the jet-$R$, showing a systematic pattern across all the $p_{T}^{jet}$ regions. This agreement with data boosts confidence in the theoretical descriptions of jet quenching and the impact of jet radius on energy loss and medium effects.
@@
Altogether, the JETSCAPE framework convincingly describes the complete evolution of the parton shower by adopting a virtuality-based multi-stage approach for energy loss. Across the entire $p_{T}^{jet}$ regions, the radius-dependent double ratio predictions from the LBT, MARTINI and AdS/CFT models exhibit distinct jet energy loss mechanisms. MARTINI, which combines both radiative and collisional energy loss processes, shows moderate dependence on jet-$R$, with a gradual rise in double ratio as larger radii recover more radiated energy. However, the model demonstrates greater sensitivity at higher $p_{T}$ region, where the increased contribution from gluon radiation leads to more significant changes in the double ratio. LBT exhibits a more consistent dependence on radius as compared to other two models, with its inclusion of medium recoil and dynamic effects from the QGP [58]. This allows LBT to more effectively recover the energy lost outside the smaller jet cones as the jet-$R$ increases. However, AdS/CFT, based on a strongly coupled medium, displays minimal radius dependence, as energy is predominantly dissipated into the medium with limited recovery within the jet cone. This difference highlights the varying assumptions regarding jet-medium interactions through all $p_{T}^{jet}$ regions, with LBT offering a more suitable framework for describing energy loss recovery.
@@
We extend our analysis by calculating the jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of $p_{T}^{jet}$ for (MATTER + LBT) model in Fig. 6. We observe a non-monotonous rise in the double ratio for $p_{T}^{jet}\leq$ 200 GeV, followed by a fall in the intermediate ${p_{T}}$ to the high ${p_{T}}$ region. The double ratio of larger radius jets is higher as compared to the smaller radius jets [59, 60]. As the jet is scattered in the medium, the final state partons from the medium that interact with the jet are also considered as the constituents of the jet, with many being emitted at large angles relative to the jet axis. As a result, increasing the jet radius enables the recovery of energy that would otherwise be lost outside the small jet-cone. Furthermore, the medium’s response to the jet, often referred to as the wake effect [61], contributes by redirecting particles back into the jet cone. This process can lead to an increment of ${R_{\mathrm{AA}}}$ for larger R as compared with smaller R-cones. Therefore, as more gluons are captured within the larger jet cone, the energy lost is partially gained as the area of the jet cone increases.
@@ III Conclusion
This work highlights the potential of the JETSCAPE framework to describe the energy loss and medium response phenomenon for broad-area jet cones and for very high transverse momentum jets.
@@
We express our gratitude to the JETSCAPE Collaboration for making the state-of-the-art framework publicly available for extensive research use. The authors would like to specifically thank the members of the JETSCAPE Collaboration, Yasuki Tachibana, Abhijit Majumder, Amit Kumar, and Chun Shen for their insightful discussion and valuable feedback. The authors would also like to thank IIT Mandi Param Himalaya computing facility under National Supercomputing Mission, DST, SPS local cluster facility, and the IIT Mandi SRIC seed grant support (Ref. No. IITM/SG/2024/01-2348).

### v5

@@ I Introduction
The extremely hot and dense conditions created at the start of the Big Bang led to what we now understand as the soup of deconfined state of the partons, the quark-gluon plasma (QGP) [1, 2, 3]. The QGP is of such great interest that, to study the properties of this state of matter, experimentalists have spent decades building advanced detectors capable of recreating this extremely dense, soupy state. The Relativistic Heavy-Ion Collider (RHIC) [4, 5] and the Large Hadron Collider (LHC) [6, 7, 8, 9] conduct heavy-ion collisions where the QGP is created for very short periods. During this time, the QGP medium has a substantial impact on the propagation and modification of the parton shower. In particular, the high-${p_{T}}$ jets produced in these heavy ion collisions undergo strong yield suppression and medium modification which are together referred to as jet quenching phenomena [10, 11, 12]. Jets modification has been studied in nucleus-nucleus collisions relative to proton-proton collisions to probe the properties of the QGP via constraints from model-to-data agreement  [13, 14, 15, 16]. The measurement of the nuclear modification factor for jets as well as inclusive jet fragmentation functions has revealed many important characteristics of the quark-gluon plasma [17, 18, 19, 20]. It provides strong confirmation of the interaction of partons with the deconfined plasma, the resulting medium modifications, and the eventual transition to hydrodynamic behavior in the medium [8]. Since there are many conclusive studies based on jet-${R_{AA}}$ [21, 22, 23], our effort here is to push the limits of the current event generators and various energy loss models for a better description of jet quenching phenomena at very high transverse momenta and broader jet cones with the multi-stage evolution of the parton shower using the Jet Energy-loss Tomography with a Statistically and Computationally Advanced Program Envelope (JETSCAPE) framework (version 3.6.5)  [24].
@@ II Results
This paper covers the collision energy of ${\sqrt{s_{\rm NN}}}$ = 5.02 TeV for the most central (0-10%) Pb-Pb collisions and includes a comparison with selected experimental data from the ATLAS and the CMS collaborations for high-${p_{T}}$ jets. In this work, the energy loss is modeled through the coupling of the MATTER and LBT frameworks. The secondary LBT model (which handles the low virtuality phase) remains the same throughout until and unless specified. Fig. 1 (top panel) shows the ${p+p}$ collision results for inclusive jet spectra at ${\sqrt{s}=5.02~{}\rm TeV}$ for ${|y_{jet}|<2.8}$, which follow the JETSCAPE PP-19 tune [50], are then compared to the experimental data from the ATLAS [22]. In this study, we have set running coupling constant ($\alpha_{\mathrm{s}}$) to 0.35 for all the calculations.
@@
These results validate and prove the MUSIC model’s reliability, which is adequate and highly effective in describing the experimental observations even at higher jet-${p_{T}}$ in the most central collisions. In contrast to the VISHNU model [52], MUSIC’s hydrodynamic evolution allows for a more accurate representation of the medium’s longitudinal expansion and viscous effects. This leads to a recurring trend of consistent predictions, indicating that MUSIC outperforms other hydrodynamic models in capturing the intricate dynamics of jet interactions within the quark-gluon plasma, especially at high energy scales reaching up to 1 TeV. The MUSIC model shows low sensitivity to coherence effects that dominate at high transverse momentum, compared to VISHNU (2+1)D. This enhances the reliability of MUSIC model for analyzing the jet-suppresions phenomenon at high-${p_{T}}$.
@@ II.1 Exploring the MARTINI and AdS/CFT models
Simulations of Pb-Pb collisions are performed using the MATTER model for both vacuum and medium showers, integrated with three distinct time-ordered parton shower models: LBT, MARTINI and AdS/CFT. The transition between these energy loss models is handled independently for each parton, with the switching parameter $Q_{\mathrm{SW}}$ fixed at 2 GeV. Both the MARTINI [30] and the AdS/CFT [31] models are developed to handle the low virtuality phase and simulate the energy loss once the parton is transferred from the MATTER. We replace the successful LBT model with the MARTINI, to carry out the simulations for the same conditions [47] and compare the jet-${R_{AA}}$ with the experimental data from the CMS [33]. Similarly, the calculations are done by replacing the LBT with AdS/CFT model and compared with the experimental data.
@@ II.2 Jet radius (${R}$) and jet-${p_{\mathrm{T}}}$ dependence of the ${R_{AA}}$
A study by the CMS collaboration [33] enlightens the rigorous comparisons of predictions from quenched jet event generators like HYDJET++ [54], PYQUEN [55] and theoretical models used to replicate relativistic heavy-ion collisions to the experimental data for the jet-${R_{\mathrm{AA}}}$. The article concludes with the impression that although most state-of-the-art models have progressed, there is still considerable uncertainty regarding large-area jets. The radius dependent charged-particle jet suppression analysis has also been explored in the recent ALICE study [56]. The comparison with JETSCAPE model predictions focuses on the lower jet-$p_{T}$ regime, providing valuable insight into the interplay between jet quenching and medium response for varying jet radii.
The large jet radius also indicates that the jet captures a substantial portion of the widely distributed momentum and energy deposited in the plasma. Since the JETSCAPE framework has pioneered the realization of a multi-stage approach for modular-based energy loss, the motivation here is to challenge the JETSCAPE model that has so far met our expectations in describing the variety of data observed and to investigate the limitations of the current models.
@@
For ${\mathrm{300~{}GeV}\leq p_{T}^{jet}\leq\mathrm{400~{}GeV}}$ interval, Fig. 5 (left panel) shows the predictions by the different combination of energy loss models. However, we note that there is only one measurement done by the CMS [33] in this ${p_{T}}$ interval, i.e., for $R$ = 0.3. All three models exhibit a consistent and nearly identical trend, i.e., (MATTER + LBT), (MATTER + AdS/CFT), and (MATTER + MARTINI), each predicting a minimal enhancement in the double ratio and showing good agreement with the data. This also indicates that the energy loss of the partons throughout the various jet cone sizes is comparable among LBT, MARTINI, and AdS/CFT.
@@
In the extreme high $p_{T}^{jet}$ region, specifically for ${\mathrm{500~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$, shown in Fig. 5 (right panel), the experimental data reveal that double ratio is less than unity for intermediate R. This may indicate that highly boosted jets, which possess more intricate jet substructures, undergo marginally more suppression. However, all the models moderately overestimate the jet-${R_{\mathrm{AA}}}$ double ratio in this high $p_{T}^{jet}$ region. Among them, the (MATTER + AdS/CFT) model provides the closest agreement, deviating only by ${10\%}$, and also harmonizes with the data within the uncertainties. In contrast, (MATTER + MARTINI) overpredicts the ratio, although it follows the same general pattern observed in other $p_{T}^{jet}$ regions.
Overall, all models predict that the double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) increases with the jet-$R$, showing a systematic pattern. However, ${R_{\mathrm{AA}}}$ visibly decreases at high momentum (${p_{T}^{jet}\geq\mathrm{500~{}GeV}}$). In the extreme high $p_{T}^{jet}$ region (${\mathrm{500~{}GeV}\leq p_{T}^{jet}\leq\mathrm{1~{}TeV}}$), the MATTER+LBT exhibits a marginal decreasing trend at larger radii but remains above 1. The MATTER+AdS/CFT closely aligns with the CMS data, displaying a subtle decrease in the double ratio for larger jet radii. Meanwhile, the MATTER+MARTINI tends to exceed the other two models following an increasing trend in the double ratio.
@@
Altogether, the JETSCAPE framework convincingly describes the complete evolution of the parton shower by adopting a virtuality-based multi-stage approach for energy loss. Across the entire $p_{T}^{jet}$ regions, the radius-dependent double ratio predictions from the LBT, MARTINI and AdS/CFT models exhibit distinct jet energy loss mechanisms. MARTINI, based on the AMY formalism, combines both radiative and collisional energy loss processes. This results in a minimal medium response where energy is absorbed by the thermal background without generating diffusion wakes. Therefore, it shows a moderate dependence on jet-$R$, with a gradual rise in double ratio as larger radii recover more radiated energy. However, the model demonstrates greater sensitivity in the higher $p_{T}$ region, where the increased contribution from gluon radiation leads to more significant changes in the double ratio. In contrast, LBT incorporates both elastic and inelastic scatterings along with a detailed treatment of thermal recoil partons and radiated gluons to capture the full jet-induced medium response. This leads to exhibit a more consistent dependence on radius compared to the other two models, with its inclusion of dynamic effects of the QGP [57]. This allows LBT to more effectively recover the energy lost outside the smaller jet cones through the medium excitation as the jet-$R$ increases. However, the AdS/CFT, which transitions to a strongly coupled energy loss mechanism characterized by continuous momentum degradation via a drag force, leads to greater suppression. As energy is predominantly dissipated into the medium with minimal recoil or recovery, this model exhibits a weak dependence on the jet radius, resulting in limited energy retention within the jet cone. Overall, these differences highlight the critical role of medium response and transport mechanisms in shaping jet quenching patterns, particularly in cone-size-dependent energy loss. They also reflect the varying assumptions regarding jet-medium interactions through all $p_{T}^{jet}$ regions, with LBT offering a more suitable framework to describe energy loss recovery.
@@
We extend our analysis by calculating the jet-${R_{\mathrm{AA}}}$ double ratio (${R^{\mathrm{R}}_{\mathrm{AA}}/R^{\mathrm{R=0.2}}_{\mathrm{AA}}}$) as a function of $p_{T}^{jet}$ for (MATTER + LBT) model in Fig. 6. We observe a non-monotonous rise in the double ratio for $p_{T}^{jet}\leq$ 200 GeV, followed by a fall in the intermediate ${p_{T}}$ to the high ${p_{T}}$ region. The double ratio of larger radius jets is higher as compared to the smaller radius jets [58, 56]. As the jet propagates through the medium, the final-state partons originating from jet–medium interactions are also included as jet constituents, with many of them emitted at large angles relative to the jet axis. As a result, increasing the jet radius enables the recovery of energy that would otherwise be lost outside the small jet-cone. Furthermore, the medium’s response to the jet, often referred to as the wake effect [59], contributes by redirecting particles back into the jet cone. This process can lead to an increment of ${R_{\mathrm{AA}}}$ for larger R as compared with smaller R-cones. Therefore, as more gluons are captured within the larger jet cone, the energy lost is partially gained as the area of the jet cone increases.
@@ III Conclusion
This work highlights the potential of the JETSCAPE framework to describe the energy loss and medium response phenomenon for broad-area jet cones and very high transverse momentum jets. While we observe a reasonable agreement between JETSCAPE predictions and experimental data, further retuning is necessary to enhance the accuracy of jet quenching modeling in heavy-ion collisions. This improvement requires a systematic approach, incorporating next-to-leading order (NLO) jet cross-section calculations in proton-proton collisions, a key improvement as NLO corrections capture higher-order partonic effects that enhance jet-medium interactions and improve agreement with data, particularly at lower $p_{T}$. Besides NLO corrections, precise tuning of key parameters, particularly the transport coefficient, $\hat{q}$ and strong coupling constant, $\alpha_{s}$ is essential. Along with these developments, further constraining the treatment of medium responses for large-radius jets could enhance energy redistribution and bring JETSCAPE predictions closer to experimental results.
@@
We express our gratitude to the JETSCAPE Collaboration for making the state-of-the-art framework publicly available for extensive research use. The authors would like to specifically thank the members of the JETSCAPE Collaboration, Yasuki Tachibana, Abhijit Majumder, Amit Kumar, and Chun Shen for their insightful discussion and valuable feedback. The authors would like to thank SPS local cluster facility, and the IIT Mandi SRIC seed grant support (Ref. No. IITM/SG/2024/01-2348). The authors acknowledge National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Himalaya’ at IIT Mandi, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.